OpenStudy (anonymous):

Factor -4d3 + 28d2 - 4d.

OpenStudy (anonymous):

-4d( d^2 - 7d +1)

OpenStudy (anonymous):

thank u how did u do that sorry im lost

OpenStudy (anonymous):

ok. let me show u step by step

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

-4d3 + 28d2 - 4d we see the common term in all the terms the common term is -4d now we take -4d out -4d3 + 28d2 - 4d. = -4d( d^2 - 7d +1)

OpenStudy (anonymous):

as -4d x d^2 = -d^3 -4d x -7d = 28d^2 -4d x 1 = -4d

OpenStudy (anonymous):

okay I understand